1) Sarah has a certain amount of money budgeted for tea and snacks during the work week, and she always aims to spend her entire budget. If she spends her entire budget on tea, she can afford 40 cups of tea. If she spends her entire budget on snacks, she can afford 8 snacks. What is the opportunity cost of a snack?

Provide your answer below:

_ cup(s) of tea

2) If a family spends its entire budget in a given time frame, the family can afford either 10 restaurant meals or 30 home meals. Assuming the family spends its entire budget on just these two goods, what is the opportunity cost of one extra restaurant meal in the time frame? (Round your answer to two decimal places.)

Provide your answer below:

_ home meals

3) Bobby is speaking to his friend and says, "this musical is going to cost me $60 when I buy the ticket." His friend corrects him and says, "actually, this concert will cost you more than $60 since you have to miss work." His friend is referring to the _________________.

Select the correct answer below:

a)economies of scale

b)budget constraint

c)opportunity cost

d)opportunity set

4) Define opportunity cost.

Select the correct answer below:

a) All possible consumption combinations of goods that someone can afford when all income is spent.

b) A cost that is made in the past and cannot be recovered.

c) All possible combinations of consumption that someone can afford given the prices of goods and the individual's income.

d) The measure of cost by what is given up in exchange for what is obtained.

5) Paul buys bus tickets and coffee each week with a budgeted amount of $6. The cost of a bus ticket is $0.60 and the cost of a cup of coffee is $1. Graph Paul's budget constraint by moving the endpoints of the line segment in the graph.

Provide your answer below:

6) Suppose a consumer has a budget for fast-food items of $36 per week and spends this money on two goods, hamburgers and pizzas. Suppose hamburgers cost ​​​​​​​$6 each and pizzas cost ​​​​​​​$9 each. Graph the budget constraint by moving the endpoints of the line segment in the graph.

Provide your answer below:
Answer #1

1. 5 cups of tea
Opportunity cost of a snack = Total cups of tea sacrificed/Total snacks consumed = 40/8 = 5 cups of tea

2. 3 home meals
Opportunity cost of one extra restaurant meal = Total home meals sacrificed/Total restaurant meals consumed = 30/10 = 3 home meals

3. c) opportunity cost
(As if one option is chosen then another option have to be given up)

4. d) The measure of cost by what is given up in exchange for what is obtained.
(Opportunity cost is the measure of what is gained by sacrificing something else)

Note: Post four MCQs at a time.